EXTENSION OF THE MEASUREMENT, ASSIGNMENT, AND FIT IN THE GROUND STATE OF THE TWO-TOP MOLECULE METHYL ACETATE

\underline{H. V. L. NGUYEN}, I. KLEINER, Laboratoire Interuniversitaire des Systèmes Atmosphériques (LISA), UMR 7583 (CNRS/Univ. Paris Est et Paris Diderot), Université de Paris Est, 61 avenue du Général de Gaulle, F-94010 Créteil cedex, France; S. SHIPMAN, Division of Natural Sciences, New College of Florida, 5800 Bay Shore Road, Sarasota, FL 34243-2109, USA; and K. KOBAYASHI, Department of Physics, Faculty of Science, Toyama University, 3190 Gofuku Toyama, Toyama, 930-8555, Japan.

New spectroscopic data were recorded for the two-top molecule methyl acetate with a chirped-pulse Fourier transform microwave spectrometer at room temperature and a jet cooled sub-millimeterwave spectrometer. More than 650 new lines with J up to 35 and K up to 15 were assigned. In total, approximately 1500 lines were fitted with 34 parameters using the program BELGI-C_{s}-2tops to a standard deviation close to the measurement error. More precise determinations of the top-top interaction and the J, K dependent parameters as well as the influence of the total rotation on the top-top interaction were carried out. The extension of the ground state spectrum of methyl acetate is a necessary step before searching for the first torsional excited transition in both torsional modes in the room temperature spectrum.